Every two weeks, Betty makes a $276 payment toward a car loan whose annual rate is 1.6%. If she will be debt-free in 5-years, how much interest will she pay on this loan?
Answer #1

Number of Periods of payment =52/2 =26
Rate =1.6%/26
PMT =276
Number of Periods =26*5 =130
PV of loan =PMT*((1-(1+r)^-n)/r) =276*((1-(1+1.6%/26)^-130)/(1.6%/26)) =34472.1315

Interest paid in 5 years or 130 Periods =PV*(1+r)^n -PV =34472.1315*(1+1.6%/26)^130-34472.1315 =2870.16